OpenStudy (anonymous):

here is a geometric series in summation notation what is the sum of the series?

OpenStudy (anonymous):

a. -122 b. 44 c. -44 d. 172 e. -84

OpenStudy (anonymous):

here first tem is 4 and then common ratio is -2 use S=a.(r^n -1)/(r-1) thus here reqd sum is 4(-128-1)/(-2-1) =4*43 =172